find the prime fractorization of 180

Answer by fcabanski(1391)

Divide by prime numbers until you've found all the prime numbers that, when multiplied together, equal 180.


It's best to begin by dividing by the lowest possible prime. It's even, so divide by 2.


180/2 = 90. So the first prime factor is 2.


90/2 = 45. Prime factors so far (2*2)


45/5 = 9. Prime factors so far (2*2*5)


9/3=3. 3 is a prime number so that gives 3*3.


Prime factorization of 180 is 2*2*3*3*5

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com